OpenStudy (anonymous):

Given the polynomial f(x)=10x^5−7x^3+6x+21 What one of the following is a possible rational root?

OpenStudy (anonymous):

@saseal can you explain this

OpenStudy (anonymous):

theres no rational root according to newton's method...

OpenStudy (anonymous):

thats not an option . these are the answer choices a. 9 / 6 b. 6 / 8 c. 5 / 20 d. 7 / 5

OpenStudy (anonymous):

try something around 1.056

OpenStudy (anonymous):

the graph of this thing touches the x axis only 1 time

OpenStudy (anonymous):

9/6 = 1.5 6/8=.75 5/20=.25 7/5=1.4

OpenStudy (anonymous):

ok try plugging in all the options into the equation and see which one ends up begin 0

OpenStudy (anonymous):

ok